# Co-occurrence is not Factual Association in Language Models

### v1

@@ 1 Introduction
Language models pretrained on large-scale text have been shown to encode a large amount of factual knowledge [1, 2] and are capable of utilizing knowledge in various reasoning scenarios [3, 4]. However, recent evidence suggest that language models could have poor sample efficiency in learning factual knowledge from text. When finetuned on simple textual demonstrations of novel facts, for example, “The capital city of Andoria is Copperton.”, even larger language models can fail to generalize the learned facts beyond simple question answering or utilize them well in reasoning [5, 6]. The success of learning factual knowledge in pretraining may simply be due to exposure to enough variations of common facts in massive corpora.
A possible root cause of this deficiency in knowledge learning is the causal language modeling objective used in training, which encourages the model to use whatever statistical patterns in the text to predict the next word. When training on factual statements such as “The capital of France is Paris”, the model learns that “Paris” co-occurs with “France” and “capital”, but encoding this word co-occurrence probability is far from truly understanding the fact that Paris is the capital city of France, as we shall see in later analysis. Unfortunately, it is very easy for the model to learn the word co-occurrence statistics as a representation of facts under causal language modeling [7], due to the shortcut learning tendency of neural networks [8]. Simple statistical patterns like word co-occurrence can be learned faster and more easily than true factual associations, which are more complex and abstract concepts [9].
@@
We release the synthetic corpus [^1] https://huggingface.co/datasets/amounts-tidings/Country-city-animals and the code [^2] https://github.com/amounts-tidings/fact_learning for the experiments in this work to facilitate further research on factual knowledge learning in language models.
@@ 2 Related work
Continual pretraining language models on new corpus is a common approach to systematically introduce new knowledge into the model. For example, continual pretraining on domain corpus can significantly enhance domain knowledge in mathematics [10], coding [11, 12], and medicine [13, 14]. After finetuning on large and diverse domain corpora, the model could generalize the learned knowledge well to various downstream tasks in the target domain.
@@
Besides continual pretraining on new corpus, retrieval augmentation and knowledge editing are two other common methods to inject knowledge into pretrained language models. Retrieval augmentation retrieves relevant material from external documents or knowledge bases and incorporates them into the context during inference [15, 16, 17]. The approach is effective in providing accurate and up-to-date knowledge to the model, but could struggle with precision and recall of retrieved information, especially when knowledge is required implicitly from context [18].
Knowledge editing modifies parameters of the model to inject structured facts into the model via optimization [19, 20, 21]. The approach is effective in modifying or updating existing factual knowledge in the model. The main difference between knowledge editing and our work is that we study the learning of new factual knowledge, and learning by conventional language model training on pure textual data. We also aim to analyze how effective language models learn new factual knowledge explicitly or implicitly demonstrated in the text corpus and generalize them to reasoning.
@@
Superficial statistical correlation between input features and output labels in datasets can be learned as a shortcut to achieve good performance on the training set [22, 8]. Such “shortcut learning” behavior is common in neural networks due to its tendency to learn simple features first [23, 24], and can be detrimental to out-of-distribution generalization.
Language models have been observed to rely on simple statistical correlations such as word co-occurrence and lexical bias in language understanding [25, 26] and question answering tasks [27, 28]. Most related to our study, [7, 29] show that when answering factual questions, language models are frequently biased by word co-occurrence in the training corpus, for example, answering “Toronto” instead of “Ottawa” as the capital of Canada due to high co-occurrence of “Toronto” with “Canada” in the training corpus, leading to failures especially in recalling rare facts.
@@
Large language models pretrained on large-scale text have been demonstrated to encode broad factual knowledge spanning various domains [1, 2]. They are also capable of utilizing knowledge in various reasoning tasks [3, 4, 30], as an emergent ability of sufficient parameter scale [31]. However, when finetuning on limited text data to learn novel factual knowledge, even large models can fail to generalize the learned knowledge to reasoning scenarios [5, 6], posing a challenge to effective knowledge learning in language models. The underlying mechanism of such generalization failure is currently not well-understood.
@@ 3 Co-occurrence is not factual association
We finetune pretrained transformer language models such as LLaMA 3 [32] and Gemma [33] with causal language modeling objective on the synthetic corpora. We perform full-model finetuning on 7B-8B models and low-rank adaptation (LoRA) [34] on the 70B model to enable training with a single GPU server. Training hyperparameters are described in Appendix B.
@@
Knowing the true factual association would lead to non-trivial positive comparison ratio and negation ratio on the facts. Having only the word co-occurrence statistics would lead to a high comparison ratio but a low negation ratio, i.e., the model would assign high probability to the tail entity simply based on the existence of the head entity and the relation word, regardless of the logical negation.
@@
Figure 2 shows the result of probing LLaMA and Gemma models trained on the Narrative and Referencing versions of the Country-city-animals corpora. The models heavily learns the co-occurrence statistics on the Narrative text, as indicated by a high comparison ratio and a near-zero negation ratio (even though each fact is paraphrased in 10 different ways). On the other hand, after finetuning on the Referencing text, the models’ behavior on the probes matches the behavior of knowing the true factual associations. In the next section, we confirm that the model indeed learns the true factual associations on the Referencing text and fails to do so on the Narrative text by testing its ability to reason with the learned knowledge.
@@
Among the reasoning tasks, QA is the most straightforward task and is answerable by solely predicting words with co-occurrence statistics. Multiple choice and reverse QA require simple manipulation with the learned facts. Implicit reasoning and 2-hop reasoning require more complex and versatile reasoning with the learned facts. Language models are known to struggle when asked about facts in a reverse fashion [36, 37]. Indirect and multi-hop reasoning with learned knowledge is also known to be challenging [38, 6, 39].
@@
Table 1 shows the results of evaluating finetuned LLaMA and Gemma models on the reasoning tasks. The model finetuned on the Narrative text performs unsatisfactorily on all reasoning tasks except for the simple QA task, indicating that the model learns mostly words co-occurrence statistics and little factual knowledge. This tendency to learn co-occurrence statistics seems independent of model size (see also Figure 2 and [7]). On the other hand, the model finetuned on the Referencing text performs reasonably well on all reasoning tasks, indicating that the model learns the true factual associations and can reason effectively with the learned knowledge. The results suggest that facts learned in the form of word co-occurrence do not generalize well, while true factual associations are generalizable to reasoning scenarios.
@@
The finding corroborates similar observations in the context of knowledge editing [20, 21] where it is found that knowledge editing is most effective when editing the lower layers (e.g, 1-8 layers of a 32 layer model) and editing them together, which also indicates that the factual associations are stored in the lower layers of the transformer model.
@@ 4 Improving factual association learning from text
It has been observed that language models struggle to learn factual knowledge that generalizes well from text [5, 6, 36]. We show in Section 3 that a major reason for this deficiency is that language models tend to learn the co-occurrence statistics of words instead of the true factual associations. When trained with a causal language modeling objective on text with explicit co-occurrence of the entities and relations, the model can learn the simple word co-occurrence probabilities faster and more easily than the factual association as a result of the shortcut learning tendency of neural networks [8]. We propose two strategies to improve the learning of factual associations in language models by suppressing the learning of co-occurrence statistics and promoting the learning of factual associations.
@@ 4.1 Learning factual associations from implicit association
The MQuAKE-T dataset [6] includes facts recently added to Wikipedia and corresponding QA questions based on the facts. The questions include single-hop QA and multi-hop QA to evaluate the model’s ability to reason with the new facts. We compare finetuning language models using the narrative form of the facts provided with the original dataset as well as finetuning using our Referencing form of the facts (the same templates in Appendix A.1 are used to generate the Referencing text as in the synthetic dataset).
@@ 4.2 Unblocking factual association learning with active forgetting
Unlike catastrophic forgetting [40, 41], where the model spontaneously forgets previously learned knowledge during finetuning, active forgetting intentionally resets parameters during training to achieve desirable learning goals. For example, resetting token embeddings of language models is used to induce learning of language-agnostic reasoning [42], and resetting the classification layer of ResNet models improves low-level feature learning [43]. Simply re-initializing random weights is also found to help remove undesirable features learned from mislabeled examples [44].
@@
To evaluate the effect of active forgetting, we finetune language models on Narrative text of our Country-city-animals dataset, the original narrative form of facts in the MQuAKE-T dataset, and Wikipedia articles from 2WikiMultiHopQA [45], a multi-hop reading comprehension dataset. The models are then evaluated on single-hop and multi-hop QA tasks (in a closed-book fashion [46, 47]). The results are shown in Table 3.
@@ 5 Conclusion
Methods for text generation: turning general text (without annotations of facts mentioned in text) into text with implicit association may require complex rewriting, for example, with the help of LLM tools such as ChatGPT [48]. Methods for rewriting general text are beyond the scope of this work.
@@
Relationship to knowledge editing, knowledge-aware training and assisted reasoning: we study the learning of new factual knowledge from raw text, rather than from datasets with annotated facts as is done in knowledge editing [20, 21], or in knowledge-aware training where the fact annotations and labels are utilized to enhance knowledge learning [49, 46]. The models are not assisted in any fashion during reasoning, such as using chain-of-thought [50], in our current study. Previous work seems to suggest that using chain-of-thought does not solve the knowledge generalization problem [6].

### v2

@@ 1 Introduction
Language models pretrained on large-scale text have been shown to encode a large amount of factual knowledge lama ; mmlu and are capable of utilizing knowledge in various reasoning scenarios bigbench ; arc . However, recent evidence suggest that language models could have poor sample efficiency in learning factual knowledge from text. When finetuned on simple textual demonstrations of novel facts, for example, “The capital city of Andoria is Copperton.”, even larger language models can fail to generalize the learned facts beyond simple question answering or utilize them well in reasoning allenzhu3.1 ; mquake . The success of learning factual knowledge in pretraining may simply be due to exposure to enough variations of common facts in massive corpora.
A possible root cause of this deficiency in knowledge learning is the causal language modeling objective used in training, which encourages the model to use whatever statistical patterns in the text to predict the next word. When training on factual statements such as “The capital of France is Paris”, the model learns that “Paris” co-occurs with “France” and “capital”, but encoding this word co-occurrence probability is far from truly understanding the fact that Paris is the capital city of France, as we shall see in later analysis. Unfortunately, it is very easy for the model to learn the word co-occurrence statistics as a representation of facts under causal language modeling knowledgecooccurrence , due to the shortcut learning tendency of neural networks shortcut . Simple statistical patterns like word co-occurrence can be learned faster and more easily than true factual associations, which are more complex and abstract concepts dissectlm .
@@
We release the synthetic corpus [^1] https://huggingface.co/datasets/xiaozeroone/Country-city-animals and the code [^2] https://github.com/xiaozeroone/fact_learning for the experiments in this work to facilitate further research on factual knowledge learning in language models.
@@ 2 Related work
Continual pretraining language models on new corpus is a common approach to systematically introduce new knowledge into the model. For example, continual pretraining on domain corpus can significantly enhance domain knowledge in mathematics (minerva, ), coding codex ; codellama , and medicine medpalm2 ; medgeminimultimodal . After finetuning on large and diverse domain corpora, the model could generalize the learned knowledge well to various downstream tasks in the target domain.
@@
Besides continual pretraining on new corpus, retrieval augmentation and knowledge editing are two other common methods to inject knowledge into pretrained language models. Retrieval augmentation retrieves relevant material from external documents or knowledge bases and incorporates them into the context during inference ragnlp ; raglm ; selfrag . The approach is effective in providing accurate and up-to-date knowledge to the model, but could struggle with precision and recall of retrieved information, especially when knowledge is required implicitly from context ragsurvey .
Knowledge editing modifies parameters of the model to inject structured facts into the model via optimization knowledgeedit ; rome ; memit . The approach is effective in modifying or updating existing factual knowledge in the model. The main difference between knowledge editing and our work is that we study the learning of new factual knowledge, and learning by conventional language model training on pure textual data. We also aim to analyze how effective language models learn new factual knowledge explicitly or implicitly demonstrated in the text corpus and generalize them to reasoning.
@@
Superficial statistical correlation between input features and output labels in datasets can be learned as a shortcut to achieve good performance on the training set databias ; shortcut . Such “shortcut learning” behavior is common in neural networks due to its tendency to learn simple features first (spectralbias, ; f-principle, ), and can be detrimental to out-of-distribution generalization.
Language models have been observed to rely on simple statistical correlations such as word co-occurrence and lexical bias in language understanding shortcutnlu ; shortcutnlu2 and question answering tasks shortcutvqa ; shortcutvqa2 . Most related to our study, knowledgecooccurrence ; knowledgemultihopcooccurrence show that when answering factual questions, language models are frequently biased by word co-occurrence in the training corpus, for example, answering “Toronto” instead of “Ottawa” as the capital of Canada due to high co-occurrence of “Toronto” with “Canada” in the training corpus, leading to failures especially in recalling rare facts.
@@
Large language models pretrained on large-scale text have been demonstrated to encode broad factual knowledge spanning various domains lama ; mmlu . They are also capable of utilizing knowledge in various reasoning tasks bigbench ; arc ; dyval , as an emergent ability of sufficient parameter scale emergent . However, when finetuning on limited text data to learn novel factual knowledge, even large models can fail to generalize the learned knowledge to reasoning scenarios allenzhu3.1 ; mquake , posing a challenge to effective knowledge learning in language models. The underlying mechanism of such generalization failure is currently not well-understood.
@@ 3 Co-occurrence is not factual association
We finetune pretrained transformer language models such as LLaMA 3 llama3 and Gemma gemma with causal language modeling objective on the synthetic corpora. We perform full-model finetuning on 7B-8B models and low-rank adaptation (LoRA) lora on the 70B model to enable training with a single GPU server. Training hyperparameters are described in Appendix B.
@@
Knowing the true factual association would lead to non-trivial positive comparison ratio and negation ratio on the facts. Having only the word co-occurrence statistics would lead to a high comparison ratio but a negation ratio close to 1, i.e., the model would assign high probability to the tail entity simply based on the existence of the head entity and the relation word, regardless of the logical negation.
@@
Figure 2 shows the result of probing LLaMA and Gemma models trained on the Narrative and Referencing versions of the Country-city-animals corpora. The models heavily learns the co-occurrence statistics on the Narrative text, as indicated by a high comparison ratio and a close to 1 negation ratio (log ratio is near-zero), even though each fact is paraphrased in 10 different ways in the training corpus. On the other hand, after finetuning on the Referencing text, the models’ behavior on the probes matches the behavior of knowing the true factual associations. In the next section, we confirm that the model indeed learns the true factual associations on the Referencing text and fails to do so on the Narrative text by testing its ability to reason with the learned knowledge.
@@
Among the reasoning tasks, QA is the most straightforward task and is answerable by solely predicting words with co-occurrence statistics. Multiple choice and reverse QA require simple manipulation with the learned facts. Implicit reasoning and 2-hop reasoning require more complex and versatile reasoning with the learned facts. Language models are known to struggle when asked about facts in a reverse fashion reversecurse ; allenzhu3.2 . Indirect and multi-hop reasoning with learned knowledge is also known to be challenging lmmultihopreason ; mquake ; lmreasonsurvey .
@@
Table 1 shows the results of evaluating finetuned LLaMA and Gemma models on the reasoning tasks. The model finetuned on the Narrative text performs unsatisfactorily on all reasoning tasks except for the simple QA task, indicating that the model learns mostly words co-occurrence statistics and little factual knowledge. This tendency to learn co-occurrence statistics seems independent of model size (see also Figure 2 and knowledgecooccurrence ). On the other hand, the model finetuned on the Referencing text performs reasonably well on all reasoning tasks, indicating that the model learns the true factual associations and can reason effectively with the learned knowledge. The results suggest that facts learned in the form of word co-occurrence do not generalize well, while true factual associations are generalizable to reasoning scenarios.
@@
The finding corroborates similar observations in the context of knowledge editing rome ; memit where it is found that knowledge editing is most effective when editing the lower layers (e.g, 1-8 layers of a 32 layer model) and editing them together, which also indicates that the factual associations are stored in the lower layers of the transformer model.
@@ 4 Improving factual association learning from text
It has been observed that language models struggle to learn factual knowledge that generalizes well from text allenzhu3.1 ; mquake ; reversecurse . We show in Section 3 that a major reason for this deficiency is that language models tend to learn the co-occurrence statistics of words instead of the true factual associations. When trained with a causal language modeling objective on text with explicit co-occurrence of the entities and relations, the model can learn the simple word co-occurrence probabilities faster and more easily than the factual association as a result of the shortcut learning tendency of neural networks shortcut . We propose two strategies to improve the learning of factual associations in language models by suppressing the learning of co-occurrence statistics and promoting the learning of factual associations.
@@ 4.1 Learning factual associations from implicit association
The MQuAKE-T dataset mquake includes facts recently added to Wikipedia and corresponding QA questions based on the facts. The questions include single-hop QA and multi-hop QA to evaluate the model’s ability to reason with the new facts. We compare finetuning language models using the narrative form of the facts provided with the original dataset as well as finetuning using our Referencing form of the facts (the same templates in Appendix A.1 are used to generate the Referencing text as in the synthetic dataset).
@@ 4.2 Unblocking factual association learning with active forgetting
Unlike catastrophic forgetting catastrophicforgetting ; catastrophicforgetting2 , where the model spontaneously forgets previously learned knowledge during finetuning, active forgetting intentionally resets parameters during training to achieve desirable learning goals. For example, resetting token embeddings of language models is used to induce learning of language-agnostic reasoning lmactiveforgetting , and resetting the classification layer of ResNet models improves low-level feature learning rifle . Simply re-initializing random weights is also found to help remove undesirable features learned from mislabeled examples activeforgetting .
@@
To evaluate the effect of active forgetting, we finetune language models on Narrative text of our Country-city-animals dataset, the original narrative form of facts in the MQuAKE-T dataset, and Wikipedia articles from 2WikiMultiHopQA 2wikimultihopqa , a multi-hop reading comprehension dataset. The models are then evaluated on single-hop and multi-hop QA tasks (in a closed-book fashion tokenmask ; conditionallm ). The results are shown in Table 3.
@@ 5 Conclusion
Methods for text generation: turning general text (without annotations of facts mentioned in text) into text with implicit association may require complex rewriting, for example, with the help of LLM tools such as ChatGPT chatgpt . Methods for rewriting general text are beyond the scope of this work.
@@
Relationship to knowledge editing, knowledge-aware training and assisted reasoning: we study the learning of new factual knowledge from raw text, rather than from datasets with annotated facts as is done in knowledge editing rome ; memit , or in knowledge-aware training where the fact annotations and labels are utilized to enhance knowledge learning knowledgeawaretraining ; tokenmask . The models are not assisted in any fashion during reasoning, such as using chain-of-thought cot , in our current study. Previous work seems to suggest that using chain-of-thought does not solve the knowledge generalization problem mquake .
Acknowledgments and Disclosure of Funding
The work is supported by Noncommunicable Chronic Diseases-National Science and Technology Major Project (Grant No. 2023ZD0506501). We thank Xien Liu and the anonymous reviewers for helpful comments and feedback.